OpenStudy (anonymous):

I need to find S25 for 1/2 +1+ 3/2 + 2+...

OpenStudy (anonymous):

if x=1/2 then, (1/2)+5(25-1)=a25? @SolomonZelman

OpenStudy (solomonzelman):

in this case it is not (1/2)+\(\large\color{red}{ \rm 5}\)(25-1) but ½+½(25-1)=a25 (when looking for 25th term of course )

OpenStudy (solomonzelman):

You already asked this question; by the way .